Reports 1-1 of 1 Clear search Modify search
Detector Operation (Operations Report)
berni - 15:56 Thursday 29 July 2021 (52694) Print this report
Operator Report - Morning shift

At the start of the shift I checked the alignment of the green (no manual alignment was needed) and then I locked the ITF in LOCKED_ARMS_IR_ALS.

Then a series of earthquakes in the Alaska region unlocked the ITF, the strongest one (magnitude 8.2, see attached picture) opened many SBE loops,  the NE IP, the NE F7; to avoid further problems I manually opened all the SBE loops and also the NE_LC.

The effect of the earthquake was present until 11:30 UTC (see attached plot 2) and prevented any kind of ISC activities.

The SBE loops have been closed around 10:30 UTC; the SBE SNEB loop was closed thanks to help of Matteo because the horizontal loop was oscillating (SBE expert have been informed).

I fully recovered the NE suspension at 11:40 UTC; I realigned the cavities on the IR and I locked the the arms; I checked again the alignment of the green laser and finally at 12:03 UTC I locked the ITF in  LOCKED_ARMS_IR_ALS.

From that point started the planned ISC activity.


Other parallel activities were on SQB1 which was vented and openend this morning.


Sub-system reports

Air Conditioning
High temperature in DER; experts verified that it is due to the presence of person in the lab.

Images attached to this report
Search Help